a rectangle is 6 feet long and has a perimeter of 20 1/3 feet what is the width
Answer by macston(5194)

L=length; W=width;P=perimeter=2(L+W)
20 1/3 ft=2(6 ft+W)
20 1/3 ft=12 ft +2W Subtract 12 from each side.
8 1/3 feet=2W Divide each side by 2
4 1/6 feet=W ANSWER: The width of the rectangle is 4 feet 2 inches.
CHECK:
P=2(L+W)
20 1/3 ft=2(6 ft+4 1/6 ft)
20 1/3 ft=2(10 1/6 ft)
20 1/3 ft=20 1/3 ft